0 3 (1 point) Find the orthogonal projection of v= -18 -14 | onto the subspace V of R³ spanned by -2.. 4 and 2. 1 projv(v) =Please write/type clearly.
Answer #1

Question Find the orthogonal projection v of subspace onto the Va 1-18 -14 -2 and -4 4 Rs spanned by 2 To find: pooj v () = ?U..v= (-4 [-4 4 2 -18 = (-4) (-18) + (4) (-14) + (2)(-2) E -14 -2 = 72 - 56 - 4 U,. V = 12 2 (0)* + (0)2 +-1) 42:42 = [o 1 -1